prove that sin 2 42 o - cos 2 78 o = (5 1/2 + 1)/8

prove that


sin242o - cos278o = (51/2 + 1)/8

Grade:11

1 Answers

vikas askiitian expert
13 years ago

LHS = sin242 - cos278

      =-(cos278 - sin242)

we have cos2A - sin2B =cos(A+B)cos(A-B)

using this we get

 LHS= -(cos120cos36)

     =(cos36)/2                                        [ cos36 = (51/2 +1)/4 ]

     =(51/2+1)/8 =RHS

hence proved